# Asymptotic and compound e-values: multiple testing and empirical Bayes

### v2

@@ 6.3 Combination and derandomization
The connections between FDR controlling procedures, e-BH, and compound e-values laid out in Sections 6.1 and 6.2 motivate the following general and practical mechanism for combining discoveries across multiple testing procedures. To be concrete, suppose we run $L$ different multiple testing procedures $\mathcal{D}_{1},\dots,\mathcal{D}_{L}$. The $\ell$-th procedure is applied to a subset of hypotheses $\{\mathcal{P}_{k}:\;k\in\mathcal{\mathcal{missing}}{S}_{\ell}\}$ where $\mathcal{S}_{\ell}\subseteq\mathcal{K}$ and controls the FDR at level $\alpha_{\ell}$. Then we may proceed as follows:
@@ 6.4 Compound e-values as weights in p-value based multiple testing
Let $P_{1},\dots,P_{K}$ be p-values for the hypotheses $H_{1},\dots,H_{K}$ and let $W_{1},\dots,W_{K}$ be $[0,\infty]$-valued. Define $Q_{k}=P_{k}/W_{k}$ with the convention $/0=0$ 0. For $k\in\mathcal{K}$, let $Q_{(k)}$ be the $k$-th order statistic of $Q_{1},\ldots,Q_{K}$, from the smallest to the largest. Consider the procedure that rejects hypothesis $k$ if $Q_{k}$ is among the smallest $k_{q}^{*}$ values $Q_{1},\dots,Q_{K}$, where
@@
with the convention $\max(\varnothing)=0$. When $W_{k}=1$ for all $k$, then the above procedure is the p-BH procedure of Benjamini and Hochberg (1995). When $W_{k}=w_{k}$ is deterministic and such that $\sum_{k\in\mathcal{K}}w_{k}=K$, then the above procedure is the weighted p-BH procedure of Genovese et al. (2006). When $W_{k}=E_{k}$ and $E_{1},\dots,E_{K}$ are compound e-values, then the above procedure is the ep-BH (e-weighted p-BH) procedure of Ignatiadis et al. (2024).
The generalization of p-BH to allow for deterministic weights by Genovese et al. (2006) was motivated by differentially prioritizing hypotheses to improve the power of p-BH while maintaining frequentist FDR control. The requirement $\sum_{k\in\mathcal{K}}w_{k}=K$ may be interpreted as a fixed size budget to be split across hypotheses. When the p-values are positive regression dependent on a subset (PRDS, Benjamini and Yekutieli (2001)) under the distributions in $\mathcal{P}$ and the weights are deterministic then existing arguments in the literature (Ramdas et al., 2019) demonstrate that the FDR of the weighted p-BH procedure is controlled at $(\alpha/K)\sum_{k\in\mathcal{N}}w_{k}$ under this assumption, which is bounded by $\alpha$ under the budget constraint $\sum_{k\in\mathcal{K}}w_{k}=K$. If the weights $W_{1},\dots,W_{K}$ are random and independent of $P_{1},\dots,P_{K}$, then by applying the above argument conditional on the weights and by iterated expectation, it follows that the FDR is controlled at $\sup_{\mathbb{P}\in\mathcal{P}}(\alpha/K)\mathbb{\sum}_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{E}^{\mathbb{P}}[W_{k}]$. The latter is controlled at $\alpha$ precisely when $W_{1},\dots,W_{K}$ are compound e-values. We note that Ignatiadis et al. (2024) defined the ep-BH procedure and proved FDR control (in their Theorem 4) when $E_{1},\dots,E_{K}$ are e-values. However, in Section 4.5, Ignatiadis et al. (2024) explained that the guarantees would continue to hold for compound e-values (although their definition is slightly different from ours).
@@ 6.5 Compound p-values in multiple testing
So far in this section we have focused on the role of compound e-values in multiple testing. Here we briefly discuss some known results about compound p-values. Armstrong (2022) provided a counterexample of jointly independent compound p-values $P_{1},\dots,P_{K}$, such that the p-BH procedure (Definition 6.6) does not control the false discovery rate.
@@
If $P_{1},\dots,P_{K}$ are (asymptotic) compound p-values, then the p-BY (Benjamini and Yekutieli, 2001) procedure (that is, the p-BH procedure applied at level $\alpha^{\prime}=\alpha\sum_{k\in\mathcal{K}}1/k$) provides (asymptotic) control of the false discovery rate at $\alpha$ under arbitrary dependence among $P_{1},\dots,P_{K}$.
@@
Second, the e-BH procedure is applied to $E_{1},\dots,E_{K}$. With this, the nonasymptotic claim then follows directly from Theorems 4.2 and 6.4. The asymptotic claim follows because calibrating asymptotic p-values yields asymptotic e-values, which when used with the e-BH procedure provides asymptotic FDR control. ∎
Furthermore, Armstrong (2022) proved that the p-BH procedure with compound p-values controls the false discovery rate under weak dependence asymptotics (as in Storey et al. (2004)); such a property of the p-BH procedure (with compound p-values) was also anticipated earlier by Efron (2007).
@@ 7 Sequence models and compound decisions
In this section we focus on sequence models, wherein the data $X$ may be written as $X=(X_{k}:k\in\mathcal{K})$ and in principle we may test each hypothesis $H_{k}$ using only $X_{k}$. We record the following definition, which we will call upon throughout this section.
@@ 7.1 Compound decision theory
7.2 The best simple separable compound e-values
@@
Suppose that $X=(X_{k}:k\in\mathcal{K})$, where $X_{k}$ takes values in a space $\mathcal{X}$. For any $\mathbb{P}\in\mathcal{P}$, let $\mathbb{P}_{k}$ be the $k$-th marginal of $\mathbb{P}$, that is, the distribution of $X_{k}$ when $(X_{k}:k\in\mathcal{K})\sim\mathbb{P}$. Suppose that for all $k\in\mathcal{K}$, the $k$-th hypothesis $H_{k}$ is specified as a simple point null hypothesis on the $k$-th marginal, i.e., $\mathcal{P}_{k}=\{\mathbb{P}\,:\,\mathbb{P}_{k}=\mathbb{P}_{k}^{\circ}\}$ for some prespecified distribution $\mathbb{P}_{k}^{\circ}$ with $\mathrm{d}\nu$-density equal to $p_{k}^{\circ}$, where $\nu$ is a common dominating measure. We assume that there exists $\mathbb{P}\in\mathcal{P}$ such that $\mathbb{P}_{k}=\mathbb{P}_{k}^{\circ}$ for all $k\in\mathcal{K}$ (as would be the case, e.g., if the $X_{k}$ are independent and we take =×$\mathbb{P}=\bigtimes_{k\in\mathcal{K}}\mathbb{P}_{k}^{\circ}$).
@@
The optimal solution to optimization problem (14) is given by the likelihood ratio of mixtures over the null and alternative:
@@
This is a well-known optimization problem (Kelly, 1956) and has solution given by (15). Finally we may verify in that $s$ in (15) indeed yields compound e-values. To this end, write $\overline{q}^{\circ}$ for the $\mathrm{d}\nu$ density of $\overline{\mathbb{Q}}^{\circ}$. Then, we have the following for any $\mathbb{P}$:
@@
  • The solution is closely related to the Optimal Discovery Procedure of Storey (2007), which solves a compound Neyman-Pearson optimization problem.
  • The objective value of the optimization problem is given by $\operatorname{KL}(\overline{\mathbb{Q}}^{\circ}\,||\,\overline{\mathbb{P}}^{\circ})$, where $\operatorname{KL}(\cdot\,||\,\cdot)$ denotes the Kullback-Leibler divergence. By convexity, $\operatorname{KL}(\overline{\mathbb{Q}}^{\circ}\,||\,\overline{\mathbb{P}}^{\circ})\leq\sum_{k\in\mathcal{K}}\operatorname{KL}(\mathbb{Q}_{k}^{\circ}\,||\,\mathbb{P}_{k}^{\circ})/K$, which means that the optimal simple separable compound e-value has a worse objective than using the optimal separable e-value for each individual testing problem of $\mathbb{P}_{k}^{\circ}$ vs $\mathbb{Q}_{k}^{\circ}$ given by $E_{k}^{*}=q_{k}^{\circ}(X_{k})/p_{k}^{\circ}(X_{k})$.
  • In view of the above, why is the optimal simple separable compound e-value of interest? The reason is that if $K$ is large (but the individual statistical problems have limited sample sizes), then it may be possible to mimic the simple separable compound e-value without knowledge of the null $\mathbb{P}_{k}^{\circ}$, e.g., via empirical Bayes methods. We believe such a strategy will be fruitful when the hypotheses $\mathcal{P}_{k}$ are composite, see Section 5.3 above and Section 7.3 below for illustrative examples of how this may work. [^5] The construction of Section 5.3 does not directly build on the compound likelihood ratio in (15) but is conceptually related.

### v3

@@ 4.3 Combination and derandomization
The connections between FDR controlling procedures, e-BH, and compound e-values laid out in Sections 4.1 and 4.2 motivate the following general and practical mechanism for combining discoveries across multiple testing procedures. To be concrete, suppose we run $L$ different multiple testing procedures $\mathcal{D}_{1},\dots,\mathcal{D}_{L}$. The $\ell$-th procedure is applied to a subset of hypotheses $\{\mathcal{P}_{k}:\;k\in\mathcal{\mathcal{missing}}{S}_{\ell}\}$ where $\mathcal{S}_{\ell}\subseteq\mathcal{K}$ and controls the FDR at level $\alpha_{\ell}$. Then we may proceed as follows:
@@
We end this subsection by using our recipe to derandomize the randomized e-BH procedures (including $R_{1}$-e-BH, $R_{2}$-e-BH, $R$-e-BH, and $U$-e-BH) proposed by Xu and Ramdas (2024). For instance, $U$-e-BH draws $U\sim\mathrm{Unif}[0,1]$ and then applies e-BH to $E_{1}/U,\ldots,E_{K}/U$. Denote any of these procedures by $\mathcal{D}$. All of these procedures have in common the following properties: they operate on e-values $E_{1},\ldots,E_{K}$ as well as an external source of randomization $U$, they control the FDR without requiring further additional assumptions over e-BH, and finally they make at least as many discoveries as e-BH (that is, $\mathcal{D}\supseteq\mathcal{D}_{\text{e-BH}}$ almost surely), often making strictly more discoveries. Both Xu and Ramdas (2024) and Lee and Ren (2024) have noted the challenge of derandomizing these procedures; naive attempts lead back to e-BH applied to the original e-values. Now consider applying the derandomization procedure of this section with $\mathcal{D}_{\ell}$ denoting an application of $\mathcal{D}$ with external randomization seed $U_{\ell}$, $S_{\ell}=\mathcal{K}$, and weights $w_{\ell}=1$. Then the resulting procedure always make at least as many discoveries as e-BH and is not identical to e-BH (i.e., it may make more discoveries).
@@ 4.4 Compound e-values as weights in p-value based multiple testing
Let $P_{1},\dots,P_{K}$ be p-variables for the hypotheses $\mathcal{P}_{1},\dots,\mathcal{P}_{K}$ and let $W_{1},\dots,W_{K}$ be $[0,\infty]$-valued. Define $Q_{k}=P_{k}/W_{k}$ with the convention $/0=0$ 0. For $k\in\mathcal{K}$, let $Q_{(k)}$ be the $k$-th order statistic of $Q_{1},\ldots,Q_{K}$, from the smallest to the largest. Consider the procedure that rejects hypothesis $k$ if $Q_{k}$ is among the smallest $k_{q}^{*}$ values $Q_{1},\dots,Q_{K}$, where
@@
with the convention $\max(\varnothing)=0$. When $W_{k}=1$ for all $k$, then the above procedure is the p-BH procedure of Benjamini and Hochberg (1995). When $W_{k}=w_{k}$ is deterministic and such that $\sum_{k\in\mathcal{K}}w_{k}=K$, then the above procedure is the weighted p-BH procedure of Genovese et al. (2006). When $W_{k}=E_{k}$ and $E_{1},\dots,E_{K}$ are compound e-variables, then the above procedure is the ep-BH (e-weighted p-BH) procedure of Ignatiadis et al. (2024).
The generalization of p-BH to allow for deterministic weights by Genovese et al. (2006) was motivated by differentially prioritizing hypotheses to improve the power of p-BH while maintaining frequentist FDR control. The requirement $\sum_{k\in\mathcal{K}}w_{k}=K$ may be interpreted as a fixed size budget to be split across hypotheses. When the p-values are positive regression dependent on a subset (PRDS, Benjamini and Yekutieli (2001)) under the distributions in $\mathcal{P}$ and the weights are deterministic then existing arguments in the literature (Ramdas et al., 2019) demonstrate that the FDR of the weighted p-BH procedure is controlled at $(\alpha/K)\sum_{k\in\mathcal{N}}w_{k}$ under this assumption, which is bounded by $\alpha$ under the budget constraint $\sum_{k\in\mathcal{K}}w_{k}=K$. If the weights $W_{1},\dots,W_{K}$ are random and independent of $P_{1},\dots,P_{K}$, then by applying the above argument conditional on the weights and by iterated expectation, it follows that the FDR is controlled at $\sup_{\mathbb{P}\in\mathcal{P}}(\alpha/K)\mathbb{\sum}_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{E}^{\mathbb{P}}[W_{k}]$. The latter is controlled at $\alpha$ precisely when $W_{1},\dots,W_{K}$ are compound e-values. We note that Ignatiadis et al. (2024) defined the ep-BH procedure and proved FDR control (in their Theorem 4) when $E_{1},\dots,E_{K}$ are e-values. However, Ignatiadis et al. (2024, Section 4.5) explained that the guarantees would continue to hold for compound e-values (although their definition is slightly different from ours).
@@ 4.5 Compound p-values in multiple testing
So far in this section we have focused on the role of compound e-values in multiple testing. Here we briefly discuss some known results about compound p-values. Armstrong (2022) provided a counterexample of jointly independent compound p-values $P_{1},\dots,P_{K}$, such that the p-BH procedure (Definition 4.6) does not control the FDR.
@@
If $P_{1},\dots,P_{K}$ are (asymptotic) compound p-values, then the p-BY (Benjamini and Yekutieli, 2001) procedure (that is, the p-BH procedure applied at level $\alpha^{\prime}=\alpha\sum_{k\in\mathcal{K}}1/k$) provides (asymptotic) control of the FDR at $\alpha$ under arbitrary dependence among $P_{1},\dots,P_{K}$.
@@
Second, the e-BH procedure is applied to $E_{1},\dots,E_{K}$. With this, the nonasymptotic claim then follows directly from Theorems 2.3 and 4.4. The asymptotic claim follows because calibrating asymptotic p-values yields asymptotic e-values, which when used with the e-BH procedure provides asymptotic FDR control. ∎
Furthermore, Armstrong (2022) proved that the p-BH procedure with compound p-values controls the FDR under weak dependence asymptotics (as in Storey et al. (2004)); such a property of the p-BH procedure (with compound p-values) was also anticipated earlier by Efron (2007).
@@ 5 Sequence models and compound decisions
In this section we focus on sequence models, wherein the data $X$ may be written as $X=(X_{k}:k\in\mathcal{K})$ and in principle we may test each hypothesis $\mathcal{P}_{k}$ using only $X_{k}$. We record the following definition, which we will call upon throughout this section.
@@ 5.1 Compound decision theory
5.2 The optimal discovery compound e-values
@@
Suppose that $X=(X_{k}:k\in\mathcal{K})$, where $X_{k}$ takes values in a space $\mathcal{X}$. For any $\mathbb{P}\in\mathcal{P}$, let $\mathbb{P}_{k}$ be the $k$-th marginal of $\mathbb{P}$, that is, the distribution of $X_{k}$ when $(X_{k}:k\in\mathcal{K})\sim\mathbb{P}$. Suppose that for all $k\in\mathcal{K}$, the $k$-th hypothesis $\mathcal{P}_{k}$ is specified as a simple point null hypothesis on the $k$-th marginal, i.e., $\mathcal{P}_{k}=\{\mathbb{P}\,:\,\mathbb{P}_{k}=\mathbb{P}_{k}^{\circ}\}$ for some prespecified distribution $\mathbb{P}_{k}^{\circ}$ with $\mathrm{d}\nu$-density equal to $p_{k}^{\circ}$, where $\nu$ is a common dominating measure. We assume that there exists $\mathbb{P}\in\mathcal{P}$ such that $\mathbb{P}_{k}=\mathbb{P}_{k}^{\circ}$ for all $k\in\mathcal{K}$ (as would be the case, e.g., if the $X_{k}$ are independent and we take the product measure $\mathbb{P}=\bigotimes_{k\in\mathcal{K}}\mathbb{P}_{k}^{\circ}$).
@@
The optimal solution to optimization problem (9) is given by the likelihood ratio of mixtures over the alternative and the null:
@@
This is a well-known optimization problem (Kelly, 1956) and has solution given by (10). Finally we may verify that $s(\cdot)$ in (10) indeed yields compound e-values. To this end, write $\overline{q}^{\circ}$ for the $\mathrm{d}\nu$ density of $\overline{\mathbb{Q}}^{\circ}$. Then, we have the following for any $\mathbb{P}$:
@@
The best simple separable compound e-values in (10) take precisely the form of the statistics of the optimal discovery procedure (ODP) of Storey (2007) and Storey, Dai, and Leek (2007). We briefly recall the statistical guarantee of the ODP. We call a multiple testing procedure $\mathcal{D}$ simple and separable (cf. Definition 5.1) if there exists a function $\phi:\mathcal{X}\to\{0,1\}$ such that $\mathds{1}_{\{k\in\mathcal{D}\}}=\phi(X_{k})$ for all $k\in\mathcal{K}$Storey (2007) proves the following.
Proposition 5.3 (Neyman-Pearson optimality of ODP).
Let $s(\cdot)$ be defined as in (10). Then, for any function $\phi:\mathcal{X}\to\{0,1\}$ and any $c>0$, the following implication holds:
$\sum_{k\in\mathcal{K}}\mathbb{E}^{\mathbb{P}_{k}}[\phi(X_{k})]\leq\sum_{k\in\mathcal{K}}\mathbb{E}^{\mathbb{P}_{k}}[\mathds{1}_{\{s(X_{k})\geq c\}}]\,\;\Longrightarrow\;\,\sum_{k\in\mathcal{K}}\mathbb{E}^{\mathbb{Q}_{k}}[\phi(X_{k})]\leq\sum_{k\in\mathcal{K}}\mathbb{E}^{\mathbb{Q}_{k}}[\mathds{1}_{\{s(X_{k})\geq c\}}].$
The interpretation of the result is as follows (Storey, 2007, Section 5.2). Suppose for each $k\in\mathcal{K}$ we draw $X_{k}$ with probability $\pi_{0}\in(0,1)$ from $\mathbb{P}_{k}$ and probability $1-\pi_{0}$ from $\mathbb{Q}_{k}$. Then the procedure that thresholds the optimal discovery statistics $s(X_{k})$ maximizes the expected number of true discoveries among all simple separable multiple testing procedures with the same expected number of false discoveries. Together with Theorem 5.2, this motivates our terminology: we call $s(X_{1}),\ldots,s(X_{k})$ the optimal discovery compound e-values
@@
  • The objective value of the optimization problem is given by $\operatorname{KL}(\overline{\mathbb{Q}}^{\circ}\,||\,\overline{\mathbb{P}}^{\circ})$, where $\operatorname{KL}(\cdot\,||\,\cdot)$ denotes the Kullback-Leibler divergence. By convexity, $\operatorname{KL}(\overline{\mathbb{Q}}^{\circ}\,||\,\overline{\mathbb{P}}^{\circ})\leq\sum_{k\in\mathcal{K}}\operatorname{KL}(\mathbb{Q}_{k}^{\circ}\,||\,\mathbb{P}_{k}^{\circ})/K$, which means that the optimal discovery compound e-values have a worse objective than using the optimal separable e-value for each individual testing problem of $\mathbb{P}_{k}^{\circ}$ vs $\mathbb{Q}_{k}^{\circ}$ given by $E_{k}^{*}=q_{k}^{\circ}(X_{k})/p_{k}^{\circ}(X_{k})$. In view of the above, why are the optimal discovery compound e-values of interest? The reason is that if $K$ is large (but the individual statistical problems have limited sample sizes), then it becomes possible to mimic the optimal discovery compound e-values without knowledge of the null $\mathbb{P}_{k}^{\circ}$, e.g., via empirical Bayes methods. Such a strategy is fruitful when the hypotheses $\mathcal{P}_{k}$ are composite. We demonstrate such a strategy explicitly in Section 6 on simultaneous t-tests where the nuisance parameters are the variances.
@@
5.3 Localized and compound universal inference
We now slightly modify the setting of Section 5.2 to accommodate composite null hypotheses with nuisance parameters. We continue to suppose that $X=(X_{k}:k\in\mathcal{K})$, where $X_{k}$ takes values in a space $\mathcal{X}$ and that the $k$-th hypothesis $\mathcal{P}_{k}$ is specified solely in terms of the marginal distribution of $X_{k}$. However, we now suppose that $\mathcal{P}_{k}$ is composite and that $k$-th null marginal may be parameterized by a nuisance parameter $\psi_{k}$ that lies in a parameter space $\Psi$. Concretely, $\mathcal{P}_{k}=\{\mathbb{P}:\mathrm{d}\mathbb{P}_{k}/\mathrm{d}\nu=p^{\circ}_{\psi_{k}}\text{ for some }\psi_{k}\in\Psi\}$, where $\{p^{\circ}_{\psi}:\psi\in\Psi\}$ is a family of $\mathrm{d}\nu$-densities on $\mathcal{X}$. We seek to test each composite null $\mathcal{P}_{k}$ against the alternative $\{\mathbb{Q}:\mathrm{d}\mathbb{Q}_{k}/\mathrm{d}\nu=q^{\circ}\}$, where $q^{\circ}$ is a further $\mathrm{d}\nu$-density on $\mathcal{X}$ (which we choose to be the same for all $k\in\mathcal{K}$).
In the above setting, universal inference (UI) of Wasserman et al. (2020) provides a flexible and practical way of constructing separable e-values: [^5] Throughout this section we assume for simplicity that infima are attained.
It is easy to see that $E_{k}^{\mathrm{UI}}$ is a valid e-value.
We now propose an alternative to the universal inference e-value above that we call localized universal inference (LUI) e-value. Let $\widehat{\Psi}_{k}(\delta)\subseteq\Psi$ be a $(1-\delta)$-confidence set for the nuisance parameter $\psi_{k}$. Then define:
$E_{k}^{\mathrm{LUI}}=\frac{q^{\circ}(X_{k})}{p^{\circ}_{\widehat{\psi}_{k}^{\mathrm{LUI}}}(X_{k})},\,\,\,\;\widehat{\psi}_{k}^{\mathrm{LUI}}\in\operatorname*{argmin}\left\{\frac{q^{\circ}(X_{k})}{p^{\circ}_{\psi}(X_{k})}\,:\,\psi\in\widehat{\Psi}_{k}(\delta)\right\}.$ (12)
We have the following result.
Proposition 5.5.
Suppose that $\widehat{\Psi}_{k}(\delta)$ is a $(1-\delta)$-confidence set for $\psi_{k}$ for all $\mathbb{P}\in\mathcal{P}_{k}$. Then $E_{k}^{\mathrm{LUI}}$ defined in (12) is a $(0,\delta)$-approximate e-variable for $\mathcal{P}_{k}$.
The construction of LUI e-values is analogous to the construction of p-values by maximizing over a confidence set for the nuisance parameter by Berger and Boos (1994). Notice that if $\widehat{\psi}_{k}^{\mathrm{UI}}\in\widehat{\Psi}_{k}(\delta)$, then $E_{k}^{\mathrm{LUI}}=E_{k}^{\mathrm{UI}}$. This simple observation allows us to highlight the difference of UI and LUI. Under the alternative, $X_{k}\sim\mathbb{Q}_{k}$, it may hold (depending on $\mathbb{Q}_{k}$ and the construction of $\widehat{\Psi}_{k}(\delta)$) that $\widehat{\psi}_{k}^{\mathrm{UI}}\notin\widehat{\Psi}_{k}(\delta)$ with high probability and thus also that $E_{k}^{\mathrm{UI}}<E_{k}^{\mathrm{LUI}}$.
Now suppose instead that $\widehat{\Psi}^{K}(\delta)\subseteq\Psi^{K}$ is a $(1-\delta)$-simultaneous confidence set for $\boldsymbol{\psi}=(\psi_{1},\ldots,\psi_{K})$ and let $\widehat{\Psi}_{k}(\delta)$ be the projection of $\widehat{\Psi}^{K}(\delta)$ onto the $k$-th coordinate. Then it follows that $E_{1}^{\mathrm{LUI}},\ldots,E_{K}^{\mathrm{LUI}}$ are $(0,\delta)$-approximate compound e-variables for $\mathcal{P}_{1},\ldots,\mathcal{P}_{K}$. This construction, however, may be very conservative and does not borrow strength across problems to yield more powerful approximate compound e-variables.
There are two ideas that lead us to a more powerful construction. First, in compound decision problems, even though it may not be possible to learn individual nuisance parameters accurately (e.g., $\boldsymbol{\psi}=(\psi_{1},\ldots,\psi_{K})$), it may be possible to accurately learn the ensemble as represented by its empirical distribution
$G(\boldsymbol{\psi})=\frac{1}{K}\sum_{k\in\mathcal{K}}\delta_{\psi_{k}}.$
Specifically, let $\mathcal{G}$ be the set of all distributions supported on $\Psi$. We say that $\widehat{\mathcal{G}}(\delta)\subseteq\mathcal{G}$ is a $(1-\delta)$-localization of $G(\boldsymbol{\psi})$ if:
$\mathbb{P}(G(\boldsymbol{\psi})\in\widehat{\mathcal{G}}(\delta))\geq 1-\delta\;\text{ for all }\;\mathbb{P}\in\mathcal{P}.$
We provide an explicit example of such a construction in Section 6.1.1. The second idea is that we can profile over the nuisance parameters appearing in the optimal discovery compound e-values of Theorem 5.2 by lifting to the space of distributions. Compound universal inference (CUI) e-values are defined as follows:666 Nguyen and Gupta (2023) use universal inference for an empirical Bayes application. By contrast, here we use ideas from empirical Bayes and universal inference to provide a new construction of compound e-values.
$E_{k}^{\mathrm{CUI}}=\frac{q^{\circ}(X_{k})}{\int p^{\circ}_{\psi}(X_{k})\mathrm{d}\widehat{G}_{k}^{\mathrm{CUI}}(\psi)},\,\,\,\;\widehat{G}_{k}^{\mathrm{CUI}}\in\operatorname*{argmin}\left\{\frac{q^{\circ}(X_{k})}{\int p^{\circ}_{\psi}(X_{k})\mathrm{d}G(\psi)}\,:\,G\in\widehat{\mathcal{G}}(\delta)\right\}.$ (13)
Proposition 5.6.
Suppose that $\widehat{G}(\delta)$ is a $(1-\delta)$-localization for $G(\boldsymbol{\psi})$ for all $\mathbb{P}\in\mathcal{P}$. Then $E_{1}^{\mathrm{CUI}},\ldots,E_{1}^{\mathrm{CUI}}$ defined in (13) are $(0,\delta)$-approximate e-variables for $\mathcal{P}_{1},\ldots,\mathcal{P}_{K}$ under $\mathcal{P}$.
6 The simultaneous t-test sequence model
In this section we provide concrete examples for data-driven implementations of localized compound universal inference, optimal discovery compound e-variables, as well as of asymptotic e-variables and asymptotic compound e-variables under asymptotics in which either the number of hypotheses or the sample size grows (or both). We do so in the stylized but practically relevant setting of multiple testing with simultaneous t-tests. We refer to Smyth (2004); Westfall et al. (2004); Finos and Salmaso (2007); Bourgon et al. (2010); Lu and Stephens (2016); Guo and Romano (2017); Ignatiadis and Huber (2021); Hoff (2022) for further works that have used this setting to demonstrate the fundamental difference between single hypothesis testing and multiple testing.
$X_{k}^{i}\,{\buildrel\text{iid}\over{\sim}\,}\mathbb{P}_{k}\;\;\text{ for }\;\;i=1,\dotsc,n.$ (14)
We seek to test the hypotheses,
$\mathcal{P}_{k}=\{\mathbb{P}\in\mathcal{P}\,:\,\mathbb{E}^{\mathbb{P}_{k}}[X_{k}^{i}]=0,\,\mathrm{var}^{\mathbb{P}_{k}}(X_{k}^{i})\in(0,\infty)\}.$ (15)
Throughout we write $\mu_{k}=\mu_{k}(\mathbb{P})=\mathbb{E}^{\mathbb{P}_{k}}(X_{k}^{i})$, $\sigma_{k}^{2}=\sigma_{k}^{2}(\mathbb{P})=\mathrm{var}^{\mathbb{P}_{k}}(X_{k}^{i})$ and $\lambda_{k}=\mu_{k}/\sigma_{k}$. We will typically keep the dependence on $\mathbb{P}$ implicit. With this notation, we seek to test whether $\mu_{k}=0$ (equivalently, $\lambda_{k}=0$). To test the above hypotheses, we start by constructing summary statistics in a group-wise manner, namely, for each $k\in\mathcal{K}$,
$\bar{X}_{k}^{(n)}=\frac{1}{n}\sum_{i=1}^{n}X^{i}_{k},\;\;\;\,S^{(n)}_{k}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(X^{i}_{k})^{2}},\;\;\;\,\hat{\sigma}_{k}^{(n)}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}\left(X^{i}_{k}-\bar{X}^{(n)}_{k}\right)^{2}}.$ (16)
We consider the testing task first under a normality assumption in Section 6.1, and then without normality in Section 6.2.
6.1 Asymptotic compound e-values under normality
We first assume that all (null and non-null) marginals $\mathcal{P}_{k}$ are normal, that is we further restrict the universe $\mathcal{P}$ to $\mathcal{P}^{\mathrm{N}}$, where
$\mathcal{P}^{\mathrm{N}}=\{\mathbb{P}\in\mathcal{P}\;:\;\mathbb{P}_{k}\,\text{is normal and }\,\underaccent{\bar}{\sigma}^{2}\leq\mathrm{var}^{\mathbb{P}_{k}}(X_{k}^{i})\leq\bar{\sigma}^{2}\;\text{ for all }\;k\},$ (17)
for known constants satisfying $0<\underaccent{\bar}{\sigma}\leq\bar{\sigma}<\infty$. For $\mathbb{P}\in\mathcal{P}^{\mathrm{N}}$, the summary statistics in (16) are sufficient. In what follows we treat $n\geq 3$ as fixed and consider asymptotics as $K\to\infty$; for this reason we omit explicit dependence on $n$, for instance, we write $\bar{X}_{k}$ instead of $\bar{X}_{k}^{(n)}$.
Our setting is encompassed by the composite null setting with nuisance parameters that we described in Section 5.3. The nuisance parameters under the null ($\psi_{k}$ in Section 5.3) are given by the variances $\sigma_{k}^{2}$. Letting $\boldsymbol{\sigma}^{2}=(\sigma_{1}^{2},\ldots,\sigma_{K}^{2})$m=, a central object is the empirical distribution of the $\sigma_{k}^{2}$, that is,
$G(\boldsymbol{\sigma}^{2})=\frac{1}{K}\sum_{k\in\mathcal{K}}\delta_{\sigma_{k}^{2}}.$ (18)
Our goal is to mimic the optimal discovery compound e-values in (10) without knowledge of $G(\boldsymbol{\sigma}^{2})$.
6.1.1 Localized compound universal inference for the simultaneous t-test
Our first approach is based on localized compound universal inference (CUI) of Section 5.3. CUI requires a confidence set of distributions containing $G(\boldsymbol{\sigma}^{2})$. We can achieve this by working with the marginal distribution of the $\hat{\sigma}_{k}^{2}$ which is ancillary for the $\mu_{k}$. Concretely, let
$f_{G}(\hat{\sigma}_{k}^{2})=\int_{0}^{\infty}p(\hat{\sigma}_{k}^{2}\mid\sigma^{2})\mathrm{d}G(\sigma^{2}),\;\;\;\,p(\hat{\sigma}_{k}^{2}\mid\sigma^{2})=\frac{\nu^{\nu/2}}{\left(\sigma^{2}\right)^{\nu/2}2^{\nu/2}\Gamma(\nu/2)}\left(\hat{\sigma}_{k}^{2}\right)^{\nu/2-1}\exp\left(-\frac{\nu\hat{\sigma}_{k}^{2}}{2\sigma^{2}}\right).$
where $\nu=n-1$ denotes the degrees of freedom. Let $\mathcal{G}$ be the class of all distributions supported on $[\underaccent{\bar}{\sigma}^{2},\bar{\sigma}^{2}]$ and consider:
$\widehat{\mathcal{G}}(\delta)=\left\{G\in\mathcal{G}\,:\,\sup_{t\in(0,\infty)}\left|\frac{1}{K}\sum_{k\in\mathcal{K}}\mathds{1}_{\{\hat{\sigma}_{k}^{2}\leq t\}}-\int_{0}^{t}f_{G}(u)\mathrm{d}u\right|\leq\sqrt{\frac{1+\log(2/\delta)}{2K}}\right\}.$ (19)
Proposition 6.1.
For any $\mathbb{P}\in\mathcal{P}^{\mathrm{N}}$ it holds that:
$\mathbb{P}(G(\boldsymbol{\sigma}^{2})\in\widehat{\mathcal{G}}(\delta))\geq 1-\delta.$